OpenStudy (anonymous):

Suppose that the functions r and s are defined for all real numbers x as follows. r(x)=2x^2 s(x)=3x Write the expressions for (r.s)(x) and (r+s)(x) and evaluate (r-s)(2) .

OpenStudy (mertsj):

\[rs(x)=6x^3\]

OpenStudy (mertsj):

\[(r+s)(x)=2x^2+3x\]

OpenStudy (mertsj):

\[(r-s)(x)=2x^2-3x\]

OpenStudy (mertsj):

\[(r-s)(2)=2(2)^2-3(2)=8-6=2\]